# Exact Consistency Under Partial Views: Graph Colorability, Capacity, and Equality in Multi-Location Encodings

### 16 Mar 2026 v3

@@ IX-A View-fiber dimensions and capacity
Proposition IX.2 (View-fiber clique sizes).
@@
Corollary IX.3 (Matroid ranks give capacity bounds).
@@
Remark. The matroid rank $t(S)$ is the “informational dimension” captured by coordinates in $S$, directly analogous to the rank function in representable matroids studied in combinatorial optimization [10] and coding theory [6]. Here $t(S)=r$ iff $S$ determines the entire state, and any admissible view containing a basis removes confusability entirely (the induced graph is edgeless). These finite-affine, coordinate-view statements show the matroid is directly applicable to the graph/capacity arc; extensions beyond this specialization are directions for future work.
@@
The binary square $\{0,1\}^{2}$ from Section V-A is an affine space over $\mathbb{F}_{2}$ with $r=2$. Each single-coordinate view has matroid rank $t(\{1\})=t(\{2\})=1$, giving fiber size $2^{2-1}=2$ and single-view capacity bound $\Theta(G_{\{i\}})\leq 1\cdot\log 2=\log 2$. The full confusability graph (the 4-cycle) achieves this bound, so the matroid rank bound is tight in this case. The matroid viewpoint shows that the 4-cycle structure arises precisely because each view captures only half the informational dimension.
X Unit-Rate Structural Interpretation
The main graph-capacity arc studies what happens once ambiguity survives and acquires nontrivial topology. This section returns to the opposite boundary case of the same model: the unit-rate corner in which that ambiguity never appears. It records two direct interpretations of that boundary case: derivation realizes the unit-rate regime, and leaving it produces the manual-update gap proved later in Section XI.
X-A Derivation as the Single-Source Structure
An encoding system has unit independent rate when $\text{DOF}(C,F)=1$. By Proposition III.6, this is exactly the regime in which every reachable state remains coherent. By Proposition III.7, any rate above $1$ admits reachable disagreement states. Thus unit rate is the unique exact-consistency boundary, and the later rate-complexity statements are its operational cost interpretation.(L: COH1-2, BND1, RAT1, RED1)
X-B Reading the Converse at Unit Rate
When ambiguity classes are trivial, no auxiliary information is needed and unit rate suffices. When a nontrivial ambiguity class survives the observation transcript, Theorem IV.3 shows that exact recovery requires a sufficiently large side-information alphabet. If that side information is unavailable but exact correctness is still demanded, additional independently accessible support is required. That is exactly what moves the system above unit rate and leads to the update-cost lower bound of Section XI. Derivation is therefore the dependence structure that preserves visible views while removing independent rate.(L: CIA3, BND2, DER2)
X-C Rate and Manual Update Cost
An encoding may expose many visible copies of a fact, but only the independent copies govern manual synchronization cost. At unit rate, one manual update changes the authoritative source and derived views follow automatically; above unit rate, each independent location must be synchronized manually. This is the cost interpretation of the threshold.(L: BND1-2)
X-D Derived Views
Definition X.1 (Derivation).
@@ X-D Derived Views
Proposition X.2 (Derivation Preserves Coherence).
@@
By Definition X.1, the value at $L_{\text{derived}}$ is determined by the value at $L_{\text{source}}$ under admissible updates. There is therefore no reachable state in which the two locations carry incompatible values for the same fact. Since the derived location has no independently writable contribution, it does not increase DOF.  $\blacksquare$
If all encodings of $F$ except one are derived from that one, then Proposition X.2 leaves exactly one independent source, hence $\text{DOF}(C,F)=1$ and exact consistency follows from Proposition III.6.(L: DER2, COH1)
@@
XI Threshold and Rate-Complexity Corollaries
The main graph-capacity arc characterizes the structure of exact failure under partial views. This section records one downstream operational consequence at the opposite boundary: independent rate governs manual synchronization cost in the same finite deterministic model.
XI-A Cost Model
Definition XI.1 (Modification Cost Model).
Let $\delta_{F}$ be a modification to fact $F$ in encoding system $C$. The effective modification complexity $M_{\mathrm{effective}}(C,\delta_{F})$ is the number of locations that must be edited manually to preserve correctness after applying $\delta_{F}$.
Only manual edits are counted. Derived locations updated automatically by the host system contribute zero effective cost.
XI-B Upper Bound at Unit Independent Rate
Theorem XI.2 (Rate-1 Upper Bound).
@@ XI-B Upper Bound at Unit Independent Rate
XI-C Lower Bound Above the Threshold
Theorem XI.3 (Above-Threshold Lower Bound).
If fact $F$ is encoded at $n$ independent locations, then
$M_{\mathrm{effective}}(C,\delta_{F})=\Omega(n).$
(L: BND2)
@@ XI-C Lower Bound Above the Threshold
Each independent location can retain an outdated value unless it is edited directly. After a source update, every one of the $n$ independent locations can therefore witness a distinct stale copy of the same fact. Any exact repair strategy must bring each such location back into agreement, and in the worst case no single manual action can repair two independently writable locations at once. Hence the effective modification complexity grows at least linearly in $n$.  $\blacksquare$
Lemma XI.4 (Information-Constrained DOF Lower Bound).
@@
Theorem IV.3 rules out exact recovery from the available tag budget. If exact correctness is nevertheless required, the architecture must supplement the insufficient side-information channel with additional independently accessible information. In the present model, that means leaving the rate-$1$ regime and paying the synchronization cost of Theorem XI.3.  $\blacksquare$
XI-D The Unbounded Gap
Theorem XI.5 (Unbounded Gap).
The ratio between the modification cost of an architecture with $n$ independent encodings and the modification cost of a rate-$1$ architecture diverges:
$\lim_{n\to\infty}\frac{M_{\mathrm{DOF}>1}(n)}{M_{\mathrm{DOF}=1}}=\infty.$
(L: BND3)
@@ XI-D The Unbounded Gap
By Theorem XI.2, a rate-$1$ architecture has constant effective modification complexity. By Theorem XI.3, an architecture with $n$ independent encodings incurs linear cost in $n$. The ratio therefore grows without bound.  $\blacksquare$
@@
XII Operational Realizability Criterion
The main graph-capacity arc characterizes what happens once ambiguity survives and must be resolved structurally. This section turns to the opposite boundary case of the same model: when can a concrete host system realize the unit-rate corner isolated earlier, and what two structural properties are needed for that purpose? In integrity language, the question is when structural integrity can be both enforced and certified by the host itself. The resulting criterion is architectural rather than a new graph-capacity theorem.
A host system realizes the rate-$1$ regime when one location is authoritative and every secondary representation is maintained as a deterministic view. Two capabilities are required:
@@ XII Operational Realizability Criterion
    Causal update propagation: updates to the authoritative source must automatically propagate to derived locations.
@@
Definition XII.1 (Verifiable Structural Integrity).
@@
XII-A Confusability and Side Information
To connect realizability to information available at verification time, fix a fact $F$ and consider the alternatives that remain compatible with the available observations. These alternatives form the same zero-error obstruction analyzed in Sections IVVIII: if the host cannot separate the surviving ambiguity class, exact verification requires additional structural side information.
Lemma XII.2 (Confusability Clique Bound).
@@ XII-A Confusability and Side Information
XII-B The Structural Timing Constraint
Definition XII.3 (Structural Fact).
@@ XII-B The Structural Timing Constraint
Theorem XII.4 (Timing Constraint for Structural Derivation).
@@
XII-C Requirement 1: Causal Update Propagation
Definition XII.5 (Causal Update Propagation).
An encoding system has causal update propagation if every update to a source location is followed by updates to all locations derived from that source with no reachable intermediate state in which source and derived locations disagree. Equivalently, propagation is part of the same admissible update event for the purposes of reachable-state analysis, rather than an asynchronous repair step with observable stale states.
@@ XII-C Requirement 1: Causal Update Propagation
Theorem XII.6 (Causal Propagation is Necessary for Unit Rate).
Achieving independent rate $1$ for structural facts requires causal update propagation.
(L: REQ2)
@@
By Theorem XII.4, structural derivation must occur at the moment the source-side structural encoding is fixed. If causal propagation is absent, then some derived location remains stale until a later manual action. During that interval the source and derived locations disagree, so the architecture does not realize exact consistency. Hence a verifiable rate-$1$ realization for structural facts requires causal propagation.  $\blacksquare$
XII-D Requirement 2: Provenance Observability
Definition XII.7 (Provenance Observability).
An encoding system has provenance observability if it supports queries that reveal which locations encode a fact, which of those locations are authoritative, and which are derived.
Provenance observability is the structural side information needed to certify that the system is genuinely in the single-source regime rather than merely appearing coherent on a small sample of states.
Theorem XII.8 (Provenance Observability is Necessary for Verifiable Unit Rate).
Verifying that independent rate $1$ holds requires provenance observability.
(L: REQ3)
@@ XII-D Requirement 2: Provenance Observability
To verify independent rate $1$, one must enumerate the locations encoding the fact, identify the authoritative source, and confirm that every remaining location is derived from it. Without provenance queries, these checks cannot be completed from inside the host system. The architecture may be coherent on observed states, but the single-source claim is not verifiable.  $\blacksquare$
XII-E Independence of the Two Requirements
@@
Theorem XII.9 (Requirements are Independent).
@@
    An encoding system can have causal propagation without provenance observability.
@@
    An encoding system can have provenance observability without causal propagation.
(L: IND1-2)
@@
XII-F The Realizability Theorem
Corollary XII.10 (Operational Realizability Criterion).
An encoding system can achieve verifiable independent rate $1$, equivalently verifiable structural integrity, for structural facts if and only if it provides both causal update propagation and provenance observability.
(L: SOT1)
@@ XII-F The Realizability Theorem
Necessity follows from Theorems XII.6 and XII.8. For sufficiency, assume both properties hold. Causal propagation ensures that every derived location is updated as part of the same structural event as the source; provenance observability then certifies that all secondary encodings are in fact derived from that source. The architecture therefore realizes a verifiable single-source encoding, i.e., independent rate $1$, which is exactly verifiable structural integrity in the sense of Definition XII.1.  $\blacksquare$
Applied to representative hosts, this criterion separates systems in which propagation and provenance are both host-native from systems missing one or both capabilities. Appendix A collects those representative readings in one table, and Supplement A contains the companion case-study traces and measurement notes.
XIII Related Work
XIII-A Zero-Error and Side-Information Lineage
Shannon’s zero-error framework and its graph-theoretic refinements by Körner and Lovász provide the closest conceptual starting point [1, 2, 3]. Witsenhausen’s zero-error side-information problem is directly relevant because it studies exact recovery under side information through graph coloring and label budgets [4]. Our one-shot coloring statements lie in that lineage. As in Witsenhausen’s formulation, an observation law induces the relevant confusability graph. The difference is that here the observation law is generated by a deterministic multi-fact tuple-space architecture with admissible coordinate-subset views, so the model itself constrains what graphs can arise.
@@ XIII-A Zero-Error and Side-Information Lineage
Both settings can be viewed as starting from an observation law and passing to the induced confusability graph, so there is no sharp “induced” versus “given” divide. The difference is architectural and structural. Witsenhausen’s setting starts from a side-information law and studies the coding problem for the graph it induces. Our setting starts from a deterministic partial-view architecture on latent tuples, with observations obtained by admissible coordinate-subset projections, and asks what graph structure that architecture can generate and what recovery laws follow from it. This makes two model-side issues central: how the admissible view family constrains the induced graph class, and how changing the view topology changes the resulting failure graph and recovery budget. In the exact full-tuple-space coordinate-view model analyzed here, that graph class can now be characterized exactly: the realizable confusability relations are precisely those determined by upward-closed families of coordinate-agreement sets. The contribution is therefore not a new coloring theorem for arbitrary graphs, but an architectural realization theory connecting multi-fact tuple-space views, a fully characterized model-generated graph class, and exact zero-error recovery.
Slepian–Wolf coding and classical entropy converses supply the second line of influence [5, 8, 9, 6]. Coding-for-computing, functional compression, and graph-entropy/characteristic-graph viewpoints are also close neighbors because they study deterministic decoder targets and coding questions once an underlying graph or function structure has been specified [11, 12, 2].
Characteristic‑graph vs. coordinate‑projection models. Characteristic‑graph and functional‑compression formalisms handle arbitrary deterministic functions of the source and therefore subsume a broader class of recovery tasks than the present model. Our model intentionally restricts attention to coordinate‑projection observations on labeled tuple spaces rather than arbitrary functions. That restriction is the source of our structural leverage: it allows a reduction of confusability to agreement‑set membership and thereby enables proofs that (a) realizable confusability relations are exactly those determined by upward‑closed agreement families, (b) coordinatewise alphabet permutations act by graph automorphisms, and (c) the induced graph class is monotone and closed under the block‑composition (strong‑product) operation. These specific structural consequences do not follow from the general characteristic‑graph viewpoint and are the reasons the paper can give deductive, architecture‑first results rather than only applying general functional‑compression theorems.
@@
[…]
XIII-B Consistency-Constrained Storage and Interaction

### 31 Mar 2026 v4

@@ VIII-A View-fiber dimensions and capacity
Proposition VIII.2footnotemark: footnotetext: : MFT86 (View-fiber clique sizes).
@@
Corollary VIII.3footnotemark: footnotetext: : FM6, AFM6-12 (Matroid ranks give capacity bounds).
@@
Remark. The matroid rank $t(S)$ is the “informational dimension” captured by coordinates in $S$, directly analogous to the rank function in representable matroids studied in combinatorial optimization [25, 26, 27] and coding theory [7]. Here $t(S)=r$ iff $S$ determines the entire state, and any admissible view containing a basis removes confusability entirely (the induced graph is edgeless). These finite-affine, coordinate-view statements show the matroid is directly applicable to the graph/capacity arc; extensions beyond this specialization are directions for future work.
@@
The binary square $\{0,1\}^{2}$ from Section IV-A is an affine space over $\mathbb{F}_{2}$ with $r=2$. Each single-coordinate view has matroid rank $t(\{1\})=t(\{2\})=1$, giving fiber size $2^{2-1}=2$ and single-view capacity bound $\Theta(G_{\{i\}})\leq 1\cdot\log 2=\log 2$. The full confusability graph (the 4-cycle) achieves this bound, so the matroid rank bound is tight in this case. The matroid viewpoint shows that the 4-cycle structure arises precisely because each view captures only half the informational dimension.
IX Unit-Rate Boundary Characterization
The preceding sections characterize the zero-error capacity, strong-power scaling laws, and upper bounds for partial views when ambiguity survives. This section turns to the operational boundary: what structural properties are necessary to remain at the zero-incoherence threshold, and what is the information-theoretic penalty for leaving it? The resulting criterion is not merely a structural curiosity. When a deterministic partial-view architecture operates above unit rate, it does not merely experience operational friction; it deterministically inherits the non-clique failure topologies and bounded exact-success rates established in Sections IVVII. Section X proves that this boundary violation strictly forces an $\Omega(n)$ manual-update lower bound, and Section XI establishes the exact realizability criteria for the zero-error regime.
IX-A Derivation as the Single-Source Structure
An encoding system has unit independent rate when $\text{DOF}(C,F)=1$. By Corollary II.7, this is exactly the regime in which every reachable state remains coherent, and any rate above $1$ admits reachable disagreement states. Thus unit rate is the unique exact-consistency boundary, and the later rate-complexity statements are its operational cost interpretation.606060: COH1-2, BND1, RAT1, RED1
IX-B Reading the Converse at Unit Rate
When ambiguity classes are trivial, no auxiliary information is needed and unit rate suffices. When a nontrivial ambiguity class survives the observation transcript, Theorem III.3 shows that exact recovery requires a sufficiently large side-information alphabet. If that side information is unavailable but exact correctness is still demanded, additional independently accessible support is required. That is exactly what moves the system above unit rate and leads to the update-cost lower bound of Section X. Derivation is therefore the dependence structure that preserves visible views while removing independent rate.616161: CIA3, BND2, DER2
IX-C Rate and Manual Update Cost
An encoding may expose many visible copies of a fact, but only the independent copies govern manual synchronization cost. At unit rate, one manual update changes the authoritative source and derived views follow automatically; above unit rate, each independent location must be synchronized manually. This is the cost interpretation of the threshold.626262: BND1-2
IX-D Derived Views
Definition IX.1 (Derivation).
@@ IX-D Derived Views
Proposition IX.2footnotemark: footnotetext: : DER2 (Derivation Preserves Coherence).
@@
By Definition IX.1, the value at $L_{\text{derived}}$ is determined by the value at $L_{\text{source}}$ under admissible updates. There is therefore no reachable state in which the two locations carry incompatible values for the same fact. Since the derived location has no independently writable contribution, it does not increase DOF.  $\blacksquare$
If all encodings of $F$ except one are derived from that one, then Proposition IX.2 leaves exactly one independent source, hence $\text{DOF}(C,F)=1$ and exact consistency follows from Corollary II.7.646464: DER2, COH1
@@
X Threshold and Rate-Complexity Corollaries
The graph-capacity arc characterizes exact failure under partial views. This section gives a complementary bound: independent rate governs the minimum synchronization cost required to restore zero-error states after source modifications. When an architecture maintains $n$ independent representations without zero-delay state synchronization, the communication cost to restore joint consistency scales as $\Omega(n)$.
X-A Cost Model
Definition X.1 (Modification Cost Model).
Let $\delta_{F}$ be a modification to fact $F$ in encoding system $C$. The synchronization cost $S(C,\delta_{F})$ is the minimum number of independent state interventions required to restore a zero-error joint state after applying $\delta_{F}$. Deterministic dependent nodes contribute zero cost, as their state is compelled by the source update.
X-B Upper Bound at Unit Independent Rate
Theorem X.2footnotemark: footnotetext: : BND1 (Rate-1 Upper Bound).
@@ X-B Upper Bound at Unit Independent Rate
X-C Lower Bound Above the Threshold
Theorem X.3footnotemark: footnotetext: : BND2 (Above-Threshold Lower Bound).
If fact $F$ is encoded at $n$ independent locations without zero-delay state synchronization, then
$S(C,\delta_{F})=\Omega(n).$
Equivalently, independent rate $n$ without deterministic coupling requires linear per-update communication cost.
@@ X-C Lower Bound Above the Threshold
Each independent location can retain stale state unless acted upon directly. After a source update, every one of the $n$ independent encodings can witness a distinct stale copy. Any protocol restoring zero-error consistency must therefore address each channel separately. No single intervention can synchronize two independently writable nodes, so the cost grows at least linearly in $n$.  $\blacksquare$
Lemma X.4footnotemark: footnotetext: : CIA4 (Information-Constrained DOF Lower Bound).
@@
Theorem III.3 rules out exact recovery from the available tag budget. If exact correctness is nevertheless required, the architecture must supplement the insufficient side-information channel with additional independently accessible information. In the present model, that means leaving the rate-$1$ regime and paying the synchronization cost of Theorem X.3.  $\blacksquare$
X-D The Unbounded Gap
Theorem X.5footnotemark: footnotetext: : BND3 (Unbounded Gap).
The ratio between the synchronization cost of an architecture with $n$ independent encodings and the synchronization cost of a rate-$1$ architecture diverges:
$\lim_{n\to\infty}\frac{S_{\mathrm{DOF}>1}(n)}{S_{\mathrm{DOF}=1}}=\infty.$
@@ X-D The Unbounded Gap
By Theorem X.2, a rate-$1$ architecture has constant synchronization cost. By Theorem X.3, an architecture with $n$ independent encodings incurs linear cost in $n$. The ratio therefore grows without bound.  $\blacksquare$
@@
XI Zero-Error Realizability
The graph-capacity arc establishes the failure topology that partial views induce. This section answers the complementary question: which channel architectures realize the verifiable zero-error regime, and which do not? The resulting realizability criterion is architectural: it identifies two necessary and sufficient structural properties. It is grounded in the same finite-obstruction and confusability theory developed above. Applied to representative deterministic partial-view channels, including programming-language runtimes, database engines, and dependency managers, the criterion yields a zero-incoherence classification that demonstrates the majority of widely deployed infrastructure operates outside the verifiable zero-error regime (Appendix A).
A host system realizes the rate-$1$ regime when one location is authoritative and every secondary representation is maintained as a deterministic view. Verifiable decoding requires two capabilities:
@@ XI Zero-Error Realizability
    Zero-delay state synchronization: updates to the authoritative source propagate to derived locations as part of the same atomic update event, with no reachable intermediate disagreement.
@@
    Structural side-information: the decoder has access to topology metadata distinguishing authoritative sources from dependent nodes.
Definition XI.1 (Verifiable Structural Integrity).
@@
XI-A Confusability and Side Information
To connect realizability to information available at verification time, fix a fact $F$ and consider the alternatives that remain compatible with the available observations. These alternatives form the same zero-error obstruction analyzed in Sections IIIVII: if the host cannot separate the surviving ambiguity class, exact verification requires additional structural side information.
Lemma XI.2footnotemark: footnotetext: : OBS5 (Confusability Clique Bound).
@@ XI-A Confusability and Side Information
XI-B The Structural Timing Constraint
Definition XI.3 (Structural Fact).
@@ XI-B The Structural Timing Constraint
Theorem XI.4footnotemark: footnotetext: : REQ1, TRI1 (Timing Constraint for Structural Derivation).
@@
XI-C Requirement 1: Causal Update Propagation
Definition XI.5 (Zero-Delay State Synchronization).
An encoding system has zero-delay state synchronization if every update to a source location is followed by updates to all locations derived from that source with no reachable intermediate state in which source and derived locations disagree. Equivalently, propagation is part of the same admissible update event for the purposes of reachable-state analysis.
@@ XI-C Requirement 1: Causal Update Propagation
Theorem XI.6footnotemark: footnotetext: : REQ2 (Zero-Delay Synchronization is Necessary for Unit Rate).
Achieving independent rate $1$ for structural facts requires zero-delay state synchronization.
@@
By Theorem XI.4, structural derivation must occur at the moment the source-side structural encoding is fixed. If zero-delay synchronization is absent, then some derived location remains stale until a later manual action. During that interval the source and derived locations disagree, so the architecture does not realize exact consistency. Hence a verifiable rate-$1$ realization for structural facts requires zero-delay synchronization.  $\blacksquare$
XI-D Requirement 2: Structural Side-Information
Definition XI.7 (Structural Side-Information).
An encoding system has structural side-information if it supports queries that reveal which locations encode a fact, which of those locations are authoritative, and which are derived.
Structural side-information is the topology metadata needed to certify that the system is genuinely in the single-source regime rather than merely appearing coherent on a small sample of states. Without structural side-information, the decoder cannot distinguish a genuine zero-error configuration from a coincidentally matching stale state.
Theorem XI.8footnotemark: footnotetext: : REQ3 (Structural Side-Information is Necessary for Verifiable Unit Rate).
Verifying that independent rate $1$ holds requires structural side-information.
@@ XI-D Requirement 2: Structural Side-Information
To verify independent rate $1$, one must enumerate the locations encoding the fact, identify the authoritative source, and confirm that every remaining location is derived from it. Without structural side-information, these checks cannot be completed from inside the host system. The architecture may be coherent on observed states, but the single-source claim is not verifiable.  $\blacksquare$
XI-E Independence of the Two Requirements
@@
Theorem XI.9footnotemark: footnotetext: : IND1-2 (Requirements are Independent).
@@
    An encoding system can have zero-delay state synchronization without structural side-information.
@@
    An encoding system can have structural side-information without zero-delay state synchronization.
@@
XI-F The Realizability Theorem
Corollary XI.10footnotemark: footnotetext: : SOT1 (Operational Realizability Criterion).
An observer can verifiably decode the single-source zero-error regime if and only if the system provides both zero-delay state synchronization and structural side-information.
@@ XI-F The Realizability Theorem
Necessity follows from Theorems XI.6 and XI.8. For sufficiency, assume both properties hold. Zero-delay synchronization ensures that every derived location is updated as part of the same structural event as the source; structural side-information then certifies that all secondary encodings are in fact derived from that source. The architecture therefore realizes a verifiable single-source encoding, i.e., independent rate $1$, which is exactly verifiable structural integrity in the sense of Definition XI.1.  $\blacksquare$
Applied to representative hosts, this criterion separates systems in which propagation and provenance are both host-native from systems missing one or both capabilities.
Table A reports the zero-incoherence classification for twelve representative channel families spanning three infrastructure domains: programming-language runtimes, database engines, and dependency managers. Of these, only two achieve verifiable rate $1$ under the criterion above. The remaining ten fail at least one of the two necessary conditions. Every entry is verified against the mechanized sufficiency conditions of Corollary XI.10. The classification demonstrates that the majority of deterministic partial-view channels underlying widely deployed infrastructure operate above the zero-incoherence threshold, where the $\Omega(n)$ synchronization cost of Theorem X.3 applies and the nontrivial confusability structure of Section IV governs exact recovery.
XII Related Work
XII-A Zero-Error and Side-Information Lineage
Shannon’s zero-error framework and its graph-theoretic refinements by Körner and Lovász provide the closest conceptual starting point [1, 2, 3]. Körner–Orlitsky’s zero-error synthesis is also a direct reference point for this lineage [4]. Witsenhausen’s zero-error side-information problem is directly relevant because it studies exact recovery under side information through graph coloring and label budgets [5]. Our one-shot coloring statements lie in that lineage. As in Witsenhausen’s formulation, an observation law induces the relevant confusability graph. The difference is that here the observation law is generated by a deterministic multi-fact tuple-space architecture with admissible coordinate-subset views, so the model itself constrains what graphs can arise.
@@ XII-A Zero-Error and Side-Information Lineage
Witsenhausen’s setting starts from a side-information law and studies the coding problem for the graph it induces. Our setting starts from a deterministic partial-view architecture on latent tuples, with observations obtained by admissible coordinate-subset projections, and asks what graph structure that architecture can generate and what recovery laws follow from it. In the exact full-tuple-space coordinate-view model analyzed here, that graph class can now be characterized exactly: the realizable confusability relations are precisely those determined by upward-closed families of coordinate-agreement sets. The contribution is therefore not a new coloring theorem for arbitrary graphs, but an architectural realization theory connecting multi-fact tuple-space views, a fully characterized model-generated graph class, and exact zero-error recovery.
Slepian–Wolf coding and classical entropy converses supply the second line of influence [6, 11, 12, 7]. Coding-for-computing, functional compression, and graph-entropy/characteristic-graph viewpoints are also close neighbors because they study deterministic decoder targets and coding questions once an underlying graph or function structure has been specified [28, 8, 29, 2, 30]. Recent zero-error function-compression work also sits in this neighborhood [31, 32].
Characteristic‑graph vs. coordinate‑projection models. Characteristic‑graph formalisms handle arbitrary deterministic functions of the source, while our model restricts attention to coordinate‑projection observations on labeled tuple spaces. That restriction enables a reduction of confusability to agreement‑set membership, yielding proofs that (a) realizable confusability relations are exactly those determined by upward‑closed agreement families, (b) coordinatewise alphabet permutations act by graph automorphisms, and (c) the induced graph class is monotone and closed under block‑composition (strong‑product). These specific structural consequences do not follow from the general characteristic‑graph viewpoint.
@@
After a confusability graph is fixed, the colorability, strong-power, Shannon-capacity, and Lovász-$\vartheta$ machinery used later is classical. The novelty here is identifying transitivity of view-generated confusability as the model-side condition that produces the cluster-graph equality subclass, together with meet-witnessing and fiber coherence as structural sufficient conditions. On the upper-bound side, this follows the same classical-to-modern arc that includes Haemers-type bounds and asymptotic-spectrum viewpoints [21, 23, 33]. The paper therefore does not claim a new theorem about arbitrary graphs, but a deterministic partial-view model that generates a fully characterized monotone agreement-set graph class on the full labeled tuple space. What it does not attempt is the broader unlabeled-isomorphism classification, or the corresponding classifications for restricted realized-state families and stochastic support-overlap models.
XII-B Consistency-Constrained Storage and Interaction